# Scalable Graph Attention‐Based Instance Selection via Mini‐Batch Sampling and Hierarchical Hashing

### v1

@@ 5.3 Component Analysis
Figure 9: Mean test effectiveness across retention rate $\rho$ for different selection strategies.
Figure 10 illustrates the impact of different normalization strategies on selection effectiveness across varying retention rates, with results averaged over ten runs using different random seeds to ensure reliability. We examine four approaches: no scaling (NS), layer weighting only (SLW), layer weighting with degree scaling (SLW+D), and layer weighting with both degree scaling and softmax normalization (SLW+D+S). Layer weighting adjusts the contribution of attention weights from different GAT layers, degree scaling normalizes node importance by the square root of their in- and out-degrees, and softmax normalizes the final importance scores to a probability distribution. At moderate retention rates (0.1, retaining 10% of instances), approaches that include degree scaling (SLW+D and SLW+D+S) outperform layer weighting alone (SLW), with nearly identical performance (0.76 effectiveness) and notable stability across training epochs. This suggests that degree scaling provides significant benefits when selecting a moderate subset of instances. At low retention rates (retaining 5% of instances), layer weighting alone (SLW) achieves slightly higher effectiveness compared to other approaches. This indicates that when selecting a smaller number of instances, the relative importance of different GAT layers becomes more critical in identifying informative instances. For highly selective rates (retaining only 1%), the combination of layer weighting and degree scaling (SLW+D) consistently outperforms other approaches (0.75 vs. 0.73 for NS). Notably, adding softmax normalization provides no additional benefit over degree scaling alone. These findings validate our methodological choice to combine layer weighting and degree scaling in GAIS. While the optimal normalization strategy varies slightly with retention rate, the results clearly show that proper scaling techniques are essential for effective selection, with softmax normalization offering little additional advantage.
Figure 10: Effect of different normalization strategies on selection effectiveness across three retention ($\rho$) rates: keeping 10%, 5%, and 1% of instances.
@@
Figure 11: Qualitative results of GAIS on different datasets. The figures illustrate the original data and the selected instances using GAIS.
@@
To thoroughly assess our method’s design, we compare different GNN architectures within GAIS (specifically the DM method), where only the GAT component is replaced while maintaining the same graph construction and IS mechanisms, as shown in Table 5. The comparable effectiveness across architectures suggests that GAIS is robust regardless of the specific GNN employed. GATv2 (Brody et al., 2022) delivers solid performance, while the Graph Transformer (Shi et al., 2021) achieves slightly better effectiveness but at substantially higher computational cost, making it less practical for large-scale applications. Despite lacking explicit attention mechanisms, models like GCN (Kipf and Welling, 2017), GIN (Xu et al., 2019), and GraphSAGE (Hamilton et al., 2018) still provide viable node importance scores through gradient-based computations, though they generally yield lower performance than attention-based approaches. This underscores the inherent advantage of attention mechanisms for IS, where the explicit attention weights provide more directly useful signals for identifying important instances than the feature representations from other GNNs.
Table 5: Comparison of different GNN architectures and our proposed improvements: HeadAmp (head-wise attention amplification) and DivAttn (diversity-promoting attention). GAIS combines both improvements.
Method $AC$ $F1$ $E$ $R$ GTT
GATv2 0.825 0.669 0.801 0.969 5.848
GCN 0.817 0.655 0.795 0.973 5.452
GIN 0.821 0.659 0.799 0.972 5.946
GraphSAGE 0.824 0.678 0.798 0.967 5.574
Graph Transformer 0.825 0.673 0.802 0.971 18.435
Base GAT 0.824 0.672 0.801 0.970 4.695
GAT + HeadAmp 0.826 0.672 0.803 0.970 4.842
GAT + DivAttn 0.828 0.673 0.804 0.970 5.198
GAIS 0.832 0.692 0.807 0.969 5.326
The progressive addition of our proposed enhancements reveals their complementary benefits. Firstly, the head-wise attention amplification improves performance by allowing the model to emphasize attention heads that identify more discriminative relationships. This targeted weighting proves particularly effective, where not all relationship patterns contribute equally to identifying informative instances. The diversity-promoting mechanism further enhances results by encouraging heads to specialize in distinct patterns, preventing redundancy in what each attention head captures. Together in the full GAIS, these components achieve the best performance with minimal increase in computation.
@@
Table 6: A comparative study of GAIS (DM variant) against state-of-the-art IS methods using the effectiveness metric. For studies that did not report this metric, we computed it manually.
@@ 5.4 Comparative Study
The comparative analysis in Table 6 demonstrates GAIS’s effectiveness against state-of-the-art IS methods. The mini-batch DM variant of GAIS achieves superior performance in 31 out of 35 datasets, with particularly significant improvements in both binary and multi-class classification problems. In large, complex datasets such as census-income (0.942 vs EGDIS’s 0.836) and coil2000 (0.939 vs LDIS’s 0.816), GAIS shows substantial improvements over existing methods. For highly discriminative datasets like skin (0.990 vs GDIS’s 0.973) and thyroid (0.960 vs ICF’s 0.912), the improvements are more modest but consistent. GAIS demonstrates particularly strong scalability on massive datasets, achieving strong performance on SUSY (0.789), covertype (0.795), and pokerhand (0.577) which contain millions of instances. Interestingly, in datasets with high class imbalance or complex feature interactions, such as yeast (0.524) and contraceptive (0.478), GAIS maintains its performance advantage over both traditional methods like CNN (0.153, 0.119) and more recent approaches like EGDIS (0.423, 0.414). While there are potential variations in experimental configurations across studies (e.g., random seeds, data splitting strategies), the consistent higher performance across such diverse datasets suggests that GAIS’s fundamental advantages — its ability to capture multi-scale relationships through graph construction and adaptively weight instance importance through attention mechanisms — enable it to better identify truly informative instances compared to both density-based methods like EGDIS and learning-based clustering approaches like CIS-BE. Unlike traditional methods that rely on fixed neighborhood metrics or static cluster assignments, GAIS’s architecture allows it to learn and leverage both local and global data patterns.
@@ 5.5 Limitations
Despite GAIS’s demonstrated effectiveness, several important limitations warrant discussion. The framework currently lacks theoretical guarantees regarding the optimality of selected instances and the preservation of crucial data characteristics, relying primarily on empirical validation. The multi-level and multi-view architectures, while effective, introduce additional hyperparameters that require tuning through computationally expensive Bayesian optimization, potentially limiting deployment in resource-constrained environments. Additionally, the attention-based importance scoring mechanism, though interpretable, is constrained by the graph structure’s locality, potentially missing important global relationships that exist beyond immediate neighborhoods. This locality constraint becomes particularly relevant in high-dimensional spaces where the curse of dimensionality can affect the quality of graph construction. The framework also assumes that the input features are sufficiently informative for meaningful graph construction, as it operates directly in the original feature space without learning more complex representations. These limitations suggest valuable directions for future research, particularly in developing theoretical foundations and more efficient hyperparameter optimization strategies.
@@ 6 Conclusions
Future research should look into developing theoretical foundations for the relationship between attention patterns and instance importance, particularly for preserving decision boundaries. Additionally, exploring adaptive strategies for configuring views and levels based on dataset characteristics could make GAIS more practical across different fields.

### v2

@@ 5.3 Component Analysis
[…]
Figure 13 illustrates the impact of different normalization strategies on selection effectiveness across varying retention rates, with results averaged over ten runs using different random seeds to ensure reliability. We examine four approaches: no scaling (NS), layer weighting only (SLW), layer weighting with degree scaling (SLW+D), and layer weighting with both degree scaling and softmax normalization (SLW+D+S). Layer weighting adjusts the contribution of attention weights from different GAT layers, degree scaling normalizes node importance by the square root of their in- and out-degrees, and softmax normalizes the final importance scores to a probability distribution. At moderate retention rates (0.1, retaining 10% of instances), approaches that include degree scaling (SLW+D and SLW+D+S) outperform layer weighting alone (SLW), with nearly identical performance (0.76 effectiveness) and notable stability across training epochs. This suggests that degree scaling provides significant benefits when selecting a moderate subset of instances. At low retention rates (retaining 5% of instances), layer weighting alone (SLW) achieves slightly higher effectiveness compared to other approaches. This indicates that when selecting a smaller number of instances, the relative importance of different GAT layers becomes more critical in identifying informative instances. For highly selective rates (retaining only 1%), the combination of layer weighting and degree scaling (SLW+D) consistently outperforms other approaches (0.75 vs. 0.73 for NS). Adding softmax normalization provides no additional benefit over degree scaling alone. These findings validate our methodological choice to combine layer weighting and degree scaling in GAIS. While the optimal normalization strategy varies slightly with retention rate, the results clearly show that proper scaling techniques are essential for effective selection, with softmax normalization offering little additional advantage.
Figure 13: Effect of different normalization strategies on selection effectiveness across three retention ($\rho$) rates: keeping 10%, 5%, and 1% of instances. Four approaches compared: no scaling (NS), layer weighting only (SLW), layer weighting with degree scaling (SLW+D), and layer weighting with degree scaling plus softmax (SLW+D+S). Degree scaling approaches (SLW+D, SLW+D+S) outperform at moderate and highly selective retention ($\rho=0.1\text{ \& }0.01$).
@@
Figure 14: Qualitative comparison of original (left) versus GAIS-selected (right) instances across four datasets using distance-based mini-batch (DM) method. Different colors represent distinct classes. GAIS maintains class distributions while achieving better separation and reducing redundancy, with selected instances preserving key structural relationships and decision boundaries.
Table 6 demonstrates the incremental value of each proposed GAT enhancement through ablation for two of our approaches: DM and ML-LSH. Head-wise attention amplification (+HeadAmp) provides consistent improvements across both methods, with particularly notable gains in F1 scores (5.2% for DM, 1.9% for ML-LSH). This supports our hypothesis that learnable scaling factors enable the model to emphasize attention heads that capture more discriminative relationships. Diversity-promoting attention (+DivAttn) shows method-dependent effects: while it substantially improves DM performance (6.2% F1 improvement), it slightly reduces ML-LSH F1 scores (-0.4%). This suggests that the multi-level structure already provides sufficient attention diversity, making explicit diversity-promoting attention less beneficial or potentially redundant. When both are combined (+Both), they consistently deliver the strongest improvements in effectiveness (1.8% for both methods). This confirms their complementary nature in identifying informative instances. The addition of cross-level attention via Nyström approximation, applicable only to ML-LSH due to its hierarchical structure, provides the most substantial gains (5.6% F1, 2.1% effectiveness). This demonstrates that global information exchange across hierarchical levels enhances the model’s ability to identify instances with varying structural importance.
Table 6: Ablation study of GAT enhancements for distance-based mini-batching (DM) and multi-level locality-sensitive hashing (ML-LSH) methods. Enhancements include head-wise attention amplification (HeadAmp), diversity-promoting attention (DivAttn), and cross-level attention via Nyström approximation (CrossLevel). Percentages in parentheses show improvement ($+$) or decline (--) relative to base GAT.
Method $AC$ $F1$ $E$ $R$
DM Base GAT 0.827 0.655 0.795 0.959
+ HeadAmp 0.835 (+1.2%) 0.673 (+5.2%) 0.802 (+1.1%) 0.958 (-0.1%)
+ DivAttn 0.836 (+1.3%) 0.682 (+6.2%) 0.804 (+1.3%) 0.960 (+0.1%)
+ Both 0.837 (+1.5%) 0.680 (+6.3%) 0.807 (+1.8%) 0.963 (+0.4%)
ML-LSH Base GAT 0.824 0.652 0.794 0.961
+ HeadAmp 0.832 (+1.3%) 0.661 (+1.9%) 0.804 (+1.5%) 0.964 (+0.3%)
+ DivAttn 0.834 (+1.3%) 0.657 (-0.4%) 0.805 (+1.5%) 0.963 (+0.2%)
+ Both 0.837 (+1.6%) 0.668 (+1.3%) 0.806 (+1.8%) 0.963 (+0.2%)
+ CrossLevel 0.838 (+1.9%) 0.676 (+5.6%) 0.809 (+2.1%) 0.964 (+0.2%)
@@
We compare different GNN architectures within GAIS (specifically the DM method) to thoroughly assess our method’s design, where only the GNN component is replaced while maintaining the same graph construction and IS mechanisms, as shown in Table 7. Traditional GNN architectures (GCN, GraphSAGE, GIN (Kipf and Welling, 2017; Hamilton et al., 2018; Xu et al., 2019)) achieve comparable effectiveness scores ($0.793-0.795$), while attention-based approaches like GATv2 show similar performance ($0.795$). The Graph Transformer achieves the highest effectiveness among the baseline GNN architectures ($0.796$). However, GAIS with our enhanced GAT architecture consistently outperforms all alternatives across all considered metrics ($AC=0.837;F1=0.680;E=0.807;R=0.963$). This supports the importance of our added enhancements which help to better identify informative instances compared to base GNN architectures. Although GNNs can provide viable node importance scores through learned representations, considering attention mechanisms explicitly provides more useful signals for selection.
Table 7: Comparison of different GNN architectures within the GAIS framework using distance-based mini-batching (DM). All methods use identical graph construction and instance selection mechanisms, with only the GNN component varied. Bold indicates best performance.
Method $AC$ $F1$ $E$ $R$
GCN 0.829 0.675 0.795 0.958
GraphSAGE 0.830 0.666 0.794 0.956
GIN 0.826 0.672 0.793 0.958
GATv2 0.827 0.666 0.795 0.962
Graph Transformer 0.827 0.657 0.796 0.963
GAIS (DM) 0.837 0.680 0.807 0.963
@@
Table 8: Comparative study of GAIS using distance-based mini-batching (DM) against state-of-the-art IS methods using the effectiveness metric. For studies that did not report this metric, we computed it manually from reported accuracy and reduction values. Bold: highest; italics: 2nd highest; underline: 3rd highest effectiveness per dataset.
@@ 5.4 Comparative Study
The comparative analysis in Table 8 demonstrates GAIS’s effectiveness against state-of-the-art IS methods. The mini-batch DM variant of GAIS achieves superior performance in 31 out of 35 datasets, with particularly significant improvements in both binary and multi-class classification problems. In large, complex datasets such as census-income (0.942 vs EGDIS’s 0.836) and coil2000 (0.939 vs LDIS’s 0.816), GAIS shows substantial improvements over existing methods. For highly discriminative datasets like skin (0.990 vs GDIS’s 0.973) and thyroid (0.960 vs ICF’s 0.912), the improvements are more modest but consistent. GAIS demonstrates particularly strong scalability on massive datasets, achieving strong performance on SUSY (0.789), covertype (0.795), and pokerhand (0.577) which contain millions of instances. In datasets with high class imbalance or complex feature interactions, such as yeast (0.524) and contraceptive (0.478), GAIS maintains its performance advantage over both traditional methods like CNN (0.153, 0.119) and more recent approaches like EGDIS (0.423, 0.414). While there are potential variations in experimental configurations across studies (e.g., random seeds, data splitting strategies), the consistent higher performance across such diverse datasets suggests that GAIS’s fundamental advantages — its ability to capture multi-scale relationships through graph construction and adaptively weight instance importance through attention mechanisms — enable it to better identify truly informative instances compared to both density-based methods like EGDIS and learning-based clustering approaches like CIS-BE. Unlike traditional methods that rely on fixed neighborhood metrics or static cluster assignments, GAIS’s architecture allows it to learn and leverage both local and global data patterns.
@@ 5.5 Limitations
Despite GAIS’s demonstrated effectiveness, several important limitations warrant discussion. The framework currently lacks theoretical guarantees regarding the optimality of selected instances and the preservation of crucial data characteristics, relying primarily on empirical validation. The multi-level and multi-view architectures, while effective, introduce additional hyperparameters that require tuning through computationally expensive Bayesian optimization, potentially limiting deployment in resource-constrained environments. Also, the attention-based importance scoring mechanism, though interpretable, is constrained by the graph structure’s locality, potentially missing important global relationships that exist beyond immediate neighborhoods. This locality constraint becomes particularly relevant in high-dimensional spaces where the curse of dimensionality can affect the quality of graph construction. The framework also assumes that the input features are sufficiently informative for meaningful graph construction, as it operates directly in the original feature space without learning more complex representations. These limitations suggest valuable directions for future research, particularly in developing theoretical foundations and more efficient hyperparameter optimization strategies.
@@ 6 Conclusions
Future research should look into developing theoretical foundations for the relationship between attention patterns and instance importance, particularly for preserving decision boundaries. Also, exploring adaptive strategies for configuring views and levels based on dataset characteristics could make GAIS more practical across different fields.